OpenStudy (anonymous):

There are 2 resistor and a battery. 8Ω and 16Ω are connected in parallel to a battery of 120V. What is the total power?

OpenStudy (anonymous):

please help

OpenStudy (kropot72):

Let the resistance of the parallel combination be R ohms. \[\large R=\frac{8\times16}{8+16}\ ohms\] The total power P can be found as follows: \[\large P=\frac{V^{2}}{R}=\frac{120^{2}}{R}=(you\ can\ calculate)\ watts\]

OpenStudy (anonymous):

I thought the equation for Power was \[P=VI\] or \[P=I ^{2}R\]?

OpenStudy (anonymous):

Thank you for answering by the way!!

OpenStudy (anonymous):

are all three equations the same?

OpenStudy (kropot72):

You're welcome. If we take your first equation \[\large P=VI\ ........(1)\] we also know from Ohm's Law that \[\large I=\frac{V}{R}\ ..........(2)\] If we plug the value of I from equation (2) into equation (1), the result is \[\large P=V \times \frac{V}{R}=\frac{V^{2}}{R}\] So we have three equations to find power.

OpenStudy (anonymous):

ohhh okay I see. One more equation. While I was doing this problem, I first solved it by doing \[P=I ^{2}R\ for each resistor and add it up together to get the total power. Then I tried to double check by answer by using this equation: P=IV since "The total power is equal to the sum of the power dissipated by the individual resistors. ." However, those value where not the same......But when I used your equation P=V^2/R , the previous statement holds true.

OpenStudy (anonymous):

*I first solved the equation by doing P=I^2R

OpenStudy (kropot72):

The equation P = VI gives the same answer as the other two equations. Perhaps your calculation needs to be checked.

OpenStudy (anonymous):

yes but P=I^2R does not ( after using this equation for each resistor and adding it up)

OpenStudy (kropot72):

If you use P = I^2R for each resistor, that is calculate the power dissipated in each resistor separately, the sum of the two values of power is the same as the power calculated by using P = VI. Using P = VI, the power dissipated in the 8 ohm resistor is given by \[\large P _{1}=(\frac{120}{8})^{2}\times8\ .........(1)\] and the power dissipated in the 16 ohm resistor is given by \[\large P _{2}=(\frac{120}{16})^{2}\times16\ ......(2)\] If you add the values of power given by (1) and (2) you will get the same answer as calculating using P = VI. Do you understand?

OpenStudy (kropot72):

Using P = I^2R ......*

OpenStudy (anonymous):

Okay, I understand. I made the mistake of using the total resistor ( for P=I^2 R) instead of the individual resistor for find current Ex: (120/5.33). Thank you soooooo much for your help!!!

OpenStudy (anonymous):

for finding*

OpenStudy (kropot72):

Glad to have been of help :)